What is the mass of NaCl in 2.5L of a 0.5M solution?

Select the correct answer below:

A) 73.05 g

B) 292.2 g

C) 46.75 g

D) 58.44 g
Answer #1

volume , V = 2.5 L

use:

number of mol,

n = Molarity * Volume

= 0.5*2.5

= 1.25 mol

Molar mass of NaCl,

MM = 1*MM(Na) + 1*MM(Cl)

= 1*22.99 + 1*35.45

= 58.44 g/mol

use:

mass of NaCl,

m = number of mol * molar mass

= 1.25 mol * 58.44 g/mol

= 73.05 g

Answer: A